# ATNPA: A Unified View of Oversmoothing Alleviation in Graph Neural Networks

### v1

@@ 3.1 GNN Message Propagation
Residual Connection
in message passing scheme can be defined in Eq. (11) where $\alpha$, $\beta$ can be hyperparamter constant but can also be learned.
@@
Dense connection
is defined in Eq. (12). Being dense, it implies that embeddings at the current layer $H^{l}$ aggregate information from all preceding layers, including $l-1,l-2,\ldots$ and so on.
@@ 3.1.2 Complex Approaches: Dynamics and Recurrence Relation
Recently, diffusion-based approaches are proposed to first model the entire graph learning process as a continuous time process (second order partial differential equation PDE or ordinary differential equation ODE) and then use different methods to discretize continuous system, leading to a nuanced recurrence relation different from traditional GNN schemes.
A general graph diffusion system (assuming a static graph) can be defined in Eq. (13), where $H$ is the learned node embedding and $l$ is the layer number from the model’s perspective or iteration number from the solver’s perspective. $H^{\prime}$ and $H^{\prime\prime}$ stand for the changing rate of $H$ (first derivative) and changing rate of $H^{\prime}$ (second derivative), respectively. A time $t$ variable acting as a continuous feature propagation corresponds to GNN feature propagation with layer $l$ increases.
@@
Discretizing Eq. (13) with different discretization schemes induces a recurrence relation similar to residual-based or dense-based connection but with a more complex structure. An example of discretization could be
@@
The principle behind the diffusion-based system is that energy preserved in the physics system while the system evolving can fit into Dirichlet energy measure and a discretization method therefore keeps Dirichlet energy from exponential decay and alleviates oversmoothing accordingly.
@@ 3.2.1 Energy Regularization
As defined in Sec 2.1, oversmoothing implies that the whole system energy is exponentially decayed to zero. Random-mask based methods provide a simple solution to alleviate the oversmoothing issue with GNN-backbone. The analysis [14] has shown that with a relatively less dense graph, GCN is less likely to suffer information loss (i.e., oversmoothing). Therefore, DropEdge [16] randomly reducing the density of the graph in the beginning naturally alleviates oversmoothing. Similarly, EGNN uses orthogonal weight initialization [15] to ensure each layer’s initial energy is upper bounded at the starting point of training. Both methods are consistent with the analysis [14] that energy is related to both propagation matrix $\texttt{aug}(A)$ and learnable weight $W$.
@@ 3.2.2 Dynamics System Modeling
Instead of regulating the energy decay using normalization or other approaches, an alternative solution is to model the process as a discretized dynamic system with explicit control on how system evolves and avoid the fixed point convergence exponentially. Accordingly, physics-inspired continuous systems have been leveraged as a starting point for constructing the new family of graph learning structures. The continuous systems equipped with Dirichlet energy are augmented with non-linearity and discretized with different discretization schemes, resulting in complex recurrence relations different from traditional residual and dense-based methods [12]. Different dynamic systems provide rich properties inheriting from their continuous form analysis that traditional GNN do not have.
@@ 3.2.3 Propagation and Transformation Decoupling
3.3 ATNPA: Unified View and Categorization
@@
The ATNPA unified view, defined from Eq. (16) to Eq. (20), outlines an abstract-level framework majority GNN methods follow, with all operators being defined in previous sections. In the following, we categorize all methods into six categorizes, and review each category in details in the succeeding subsections.
3.3.1 Categorization
@@
Following the three major themes in Sec 3.2, we categorize existing alleviation methods based on critical changes they made compared with the vanilla GNN scheme, and link them to the ATNPA unified view framework.
Residual-based:
@@
Residual-based models explicitly add skip- or residual-connection to the ATNPA’s Propagation step at Eq. (19). Examples include APPNP [33], ResGCN [32], GCNII, GEN [34], EGNN, G2-gating, etc). Initial works, such as ResGCN, GCNII, are inspired by residual connection in the computer vision field [44]. Later, EGNN and G2-gating focus on improving the coefficient of each residual component under the principle of preserving Dirichlet energy among layers.
Dense-based:
@@
Dense-based methods explicitly aggregate all layer embeddings into the final embeddings, which is reflected in ATNPA’s Aggregation step at Eq. (20). Examples include JKnet [25], DCGCN [37], MixHop [38], Scattering GCN [45] and DAGNN [36].
Random-mask based:
Random-mask based models randomly mask or drop edges/nodes of the original graph, corresponding to changes in ATNPA’s Augmentation step at Eq. (16), and then use resulted stochastic graph for propagation. Examples include DropEdge, Drop-connect [39], DropMessage [40].
Energy-based:
@@
Energy-based models introduce normalization techniques that control Dirichlet energy or feature variance of the learned embeddings to explicitly optimizing the measure of oversmoothing and alleviate oversmoothing. This is reflected at ATNPA’s Normalization step at Eq. (18) and Propagation step at Eq. (19). Examples include EGNN, PairNorm, NodeNorm [24], GroupNorm [35], G2-gating.
Diffusion-based:
@@
Diffusion-based methods first model a continuous ODE or PDE related to graph diffusion equation and then discretize the continuous equation with different discretization methods. This leads to nuanced recurrence relation and possibly a combined propagation matrix learnt from both features and topology, which corresponds to ATNPA’s Augmentation step at Eq. (16) and Propagation step at Eq. (19). Examples include GraphCon, GRAND [27], ACMP [41], Neural Sheaf Diffusion [42].
Transformer-based:
@@
3.4 Residual-based Methods
@@
where $\alpha+\beta=c_{min}$ and $c_{min}$ is a positive hyperparamter chosen to satify the lower bound of initial Dirichlet energy. Its main motivation is to choose an appropriate $c_{min}$ which is an lower bound of initial Dirichlet energy to keep the Dirichlet energy in a controllable range during propagation. Therefore, EGNN is an explicit energy preserving technique compared with implicit energy control by diffusion-based methods.
@@
$H^{l}\leftarrow H^{l-1}+F_{\theta}(H^{l-1},A)$ (27)
@@
Discussion(Residual Connection):
Note that all above methods fit into ATNPA’s unified view by making changes to the Normalization (Eq. 18) and Propagation (Eq. 19) steps. In general, residual based methods have been primarily focused on learning coefficients for each component (i.e., coefficients for $H^{l}$, $H^{l-1}$, or $F_{\theta}(H^{l-1},A)$ etc.). Nevertheless, there is insufficient study and theoretical analysis about the order of each residual components in terms of their position w.r.t activation function $\sigma(\cdot)$. To date, GEN is the only work that empirically validated that order they proposed works better than GCNII.
3.5 Dense-based Methods
@@
It can be observed that the two dense-based methods share similar final aggregation scheme (i.e., final embedding can be considered as a linear combination of layer embeddings). Yet, the embedding learnt in the intermediate layers is different. JKnetcat still includes learnable parameters in the middle and keep non-linearity while DAGNN removes both parts. Without non-linearity and learnable paramters, DAGNN essentially becomes a diffusion kernel based method similar to [50]. We can see that both methods fit into ATNPA’s unified view in Transformation (Eq. 17) and Aggregation (Eq. 20), which is the key component for dense connection based method.
@@
which is still a variant of ATNPA’s Aggregation (Eq. 20) with a slight difference in the order of aggregation before convolution instead of after convolution.
Discussion(Dense Connection):
@@
3.6 Random-mask based Methods
@@
where $\texttt{Bern}(p)$ is a matrix filled with Bernoulli distribution elements and $p$ controls the drop rate. The motivation behind edge dropping is the subspace theorem [14] that indicates less connected graph leading to slow convergence of oversmoothing state. The random-mask modification fits into ATNPA’s Augmentation step (Eq. 16).
@@
3.7 Energy-based Methods
Energy-based methods share common motivation of controlling generated embeddings in each layer with constraints on either preserving Dirichlet energy or reducing feature variance. Examples of preserving Dirichlet energy include EGNN, G2-gating, PairNorm, GroupNorm, while NodeNorm reduce feature variance.
@@ 3.7 Energy-based Methods
Discussion(Normalization):
We note that both normalization and coefficient computation approaches fit into ATNPA’s unified view in Normalization (Eq. 18) and Propagation (Eq. 19). Direct normalization on features such as mean substraction and variance shifting empirically reduce model capacity and expressive power while coefficients learning seem to be a more promising direction to not only keep Dirichlet energy but also preserve model expressive power.
3.8 Diffusion-based Methods
Diffusion-based methods consider GNN learning as a continuous system and derive solutions by formulating the system’s evolving as a model propagation process. In this context, the time component $t$ in continuous system corresponds to GNN based model’s layer concept. Different discretization methods provide a complex family of methods indicated by a continuous system and most GNN based backbone can be considered as an explicit Euler discretization (only considering the recurrence relation or the Update function in GNN framework) [27]. Examples of continuous systems include GRAND, GraphCon, ACMP, Neural Sheaf Diffusion, and G2-gating (It was first reviewed as GNN backbones, but is also related to the continuous system).
GRAND [27] leverages graph diffusion PDE equations as the continuous system and performs both explicit and implicit Euler discretization. The diffusivity is modeled with an attention structure (Eq. 33) related to node features and edges (Eq. 34):
@@ 3.8 Diffusion-based Methods
$Y^{\prime}\leftarrow F_{\theta}(A,H,t)-\gamma H-\alpha Y$ (35)
$H^{\prime}\leftarrow Y$ (36)
where $H^{\prime}$ is the first order derivative with respect to time $t$ (a default setting at physics) and $Y^{\prime}$ is equivalent to $H^{\prime\prime}$. After discretization, $t$ is essentially equivalent to layer $l$ in GNN backbones. Following IMEX (implicit-explicit) time discretization [51], GraphCon obtains a new recurrence:
$Y^{n}\leftarrow Y^{n-1}$ $+\Delta(t)(\sigma(F_{\theta}(A,H^{n-1},t^{n-1}))$
$-\gamma H^{n-1}-\alpha Y^{n-1})$ (37)
$H^{n}\leftarrow H^{n-1}$ $+\Delta(t)Y^{n}$ (38)
@@
Discussion:(Continuous System)
Diffusion systems above share a common point of establishing a connection between the feature changing rate $H^{\prime}$ and the graph gradient $\sum_{j\in Neighbor(i)}|h_{i}-h_{j}|$ which is Dirichlet energy for one node. A discretization of the system then provides a unique complex recurrence relation that preserves established connections. The niche of diffusion-based methods stem from the design that the system preserves Dirichlet energy (mitigates oversmoothing) through the complex residual recurrence structure, avoiding fixed point convergence at exponential rate and small perturbation deviates the fixed point away in the GraphCon case. This makes diffusion-based method unique, compared with other works that preserve energy through explicit feature value control or coefficient control.

### v2

@@ 3.1 GNN Message Propagation
To preserve long-term information during deep layer propagation, two types of connections are commonly used in existing works, namely residual connection and dense connection:
Residual Connection in message passing scheme can be defined in Eq. (9) where $\alpha$, $\beta$ can be hyperparamter constant but can also be learned.
@@
We note that Eq. (9) provides a choice for node embeddings update to preserve part of its original information $H^{l-1}$ and $H^{1}$ rather than entirely change to its aggregated message $F_{\theta}(H^{l-1},A)$, which typically leads to smooth update.
Dense Connection is defined in Eq. (10). Being dense, it implies that embeddings at the current layer $H^{l}$ aggregate information from all preceding layers, including $l-1,l-2,\ldots$ and so on.
@@ 3.1.2 Complex Approaches: Dynamics and Recurrence Relation
Recently, physics-informed approaches are proposed to first model the entire graph learning process as a continuous time process (second order partial differential equation PDE or ordinary differential equation ODE) and then use different methods to discretize continuous system, leading to a nuanced recurrence relation different from traditional GNN schemes.
A general physics-informed system (assuming a static graph) can be defined in Eq. (11), where $H$ is the learned node embedding and $l$ is the layer number from the model’s perspective or iteration number from the solver’s perspective. $H^{\prime}$ and $H^{\prime\prime}$ stand for the changing rate of $H$ (first derivative) and changing rate of $H^{\prime}$ (second derivative), respectively. A time $t$ variable acting as a continuous feature propagation corresponds to GNN feature propagation with layer $l$ increases.
@@
Discretizaing Eq. (11) with different discretization schemes induces a recurrence relation similar to residual connection or dense connection but with a more complex structure. An example of discretization could be
@@
where Eq. (3.1.2) and Eq. (13) shows a common form with iterative form for both its first-derivative (embedding update rate) update equation and embedding update equation.
The principle behind the physics-informed system is that energy preserved in the physics system while the system evolving can fit into Dirichlet energy measure and a discretization method therefore keeps Dirichlet energy from exponential decay and alleviates oversmoothing accordingly.
@@
Figure 3: The proposed taxonomy of oversmooth alleviation approaches for deep graph neural networks. The taxonomy includes three themes and six categories of methods for oversmooth alleviation. Some approaches, such as physics-inform methods, are used by all three themes. Others, such as residual connection, dense concatenation, graph transformer, are only used by a unique theme.
@@ 3.2.1 Energy Regularization
As defined in Sec 2.1, oversmoothing implies that the whole system energy is exponentially decayed to zero. Stochastic masking based methods provide a simple solution to alleviate the oversmoothing issue with GNN-backbone. The analysis [19] has shown that with a relatively less dense graph, GCN is less likely to suffer from information loss (i.e., oversmoothing). Therefore, DropEdge [23] randomly reducing the density of the graph in the beginning naturally alleviates oversmoothing. Similarly, EGNN uses orthogonal weight initialization [26] to ensure each layer’s initial energy is upper bounded at the starting point of training. Both methods are consistent with the analysis [19] that energy is related to both propagation matrix $\texttt{aug}(A)$ and learnable weight $W$.
@@
Methods that prevent energy from decaying exponentially can be divided into three main types: structure preserved, feature normalization preserved, and coefficient preserved. The structure preserved method mostly comes from different combinations of residual connection, dense connection, and discretization of the diffusion equation, leading to different recurrence relations of the Updating function. Feature normalization preserved method includes those normalization techniques that directly apply to features such as PairNorm, NodeNorm, etc. The coefficient preserved type is about how to select the coefficient of each residual component to preserve Dirichlet energy, including EGNN, G2-gating, etc.
@@ 3.2.2 Dynamics System Modeling
Instead of regulating the energy decay using normalization or other approaches, an alternative solution is to model the process as a discretized dynamic system with explicit control on system evolving and avoidance of the fixed point convergence. To this end, physics-inspired continuous systems have been leveraged as a starting point for constructing the new family of graph learning structures. The continuous systems equipped with Dirichlet energy are augmented with non-linearity and discretized with different discretization schemes, resulting in complex recurrence relations different from traditional residual and dense-based methods [9]. Different dynamic systems provide rich properties inheriting from their continuous form analysis that traditional GNN do not have.
@@ 3.2.3 Propagation and Transformation Decoupling
3.3 GNN Oversmoothing Alleviation Taxonomy
Based on the three themes to tackle the oversmoothing in the GNNs, we propose a taxonomy in Fig. 3, which categories all methods into three themes and six subgroups. Under the newly proposed taxonomy, some approaches, such as physics-inform methods, are used by all three themes. Whereas others, such as residual connection, dense concatenation, graph transformer, are only used by a unique theme. The taxonomy helps lay the foundation for ATNPA, the unified view and categorization (Sec 4). In the next session, we will present unified view and categorization, and review representative methods in each category, including their key steps and relation to the ATNPA, as well as their rationality in tackling the oversmoothing challenges.
4 ATNPA: Unified View and Categorization
@@
The ATNPA unified view, defined from Eq. (14) to Eq. (18), outlines an abstract-level framework majority GNN methods follow, with all operators being defined in previous sections. Notice that after augmentation as in Eq. (14), it is rare to use the original matrix $A$ unless it is in a self-supervised learning settings with multi-views of graphs utilized [50]. In the following, we categorize all methods into six categorizes, and review each category in details in the succeeding subsections.
4.1 Categorization
@@
Following the three major themes in Sec 3.2, we categorize existing alleviation methods based on critical changes they made compared with the vanilla GNN scheme, and link them to the ATNPA unified view framework. First, we summarize its general properties and show their implicit connections. Then, we discuss specific methods within each category in detail.
Energy Regularization is one major theme we categorize for alleviating oversmoothing. Both Residual connection and Dense concatenation belong to this major theme with its unique property of alleviating energy decay thorough layer updates and they separately focus on adapting the propagation step Eq. (17) and aggregation step Eq. (18). Notice that some of the methods in this category also belongs to dynamic system modeling as the entire layer stacking process can be modeled as a dynamic system and alleviating oversmoothing can be considered as controlling the energy change in the system.
Residual connection:
@@
Residual-based models explicitly add skip- or residual-connection to the ATNPA’s Propagation step at Eq. (17). Examples include APPNP [36], ResGCN [35], GCNII, GEN [37], EGNN, G2-gating, GRN, etc). Initial works, such as ResGCN, APPNP are inspired by residual connection in the computer vision field [51] and only contains either skip-connection $H^{1}$ or residual-connection $H^{l-1}$. GCNII covers both skip-connection $H^{1}$ and residual connection $H^{l-1}$ to its update() function and shows better performance. However, its residual coefficient remains hyperparameter. Later, EGNN and G2-gating focus on selection of each residual coefficient under the principle of preserving Dirichlet energy among layers and generalize the residual connection trick to the gating mechanism. GRN further generalizes the constraint of Dirichlet energy to a learnable metric for better adaptation.
Dense Concatenation:
@@
Dense concatenation based methods explicitly aggregate all layer embeddings into the final embeddings, which is reflected in ATNPA’s Aggregation step at Eq. (18). Examples include JKnet [22], DenseGCN [41], MixHop [42], Scattering GCN [52] and DAGNN [40].
In addition to the residual-based energy change regularization, many methods focus on explicit energy control directly over the node embeddings learned and this corresponds to our Normalization step at Eq. (16) and Propagation step at Eq. (17).
Energy Control:
@@
Energy-controlled models introduce normalization techniques that control Dirichlet energy or feature variance of the learned embeddings to explicitly optimizing the measure of oversmoothing and alleviate oversmoothing. This is reflected at ATNPA’s Normalization step at Eq. (16) and Propagation step at Eq. (17). Examples include EGNN, PairNorm, NodeNorm [21], GroupNorm [38], G2-gating.
While the two main branches above focus on regularizing the energy change thorough layer stacking, another branch of alleviating oversmoothing for energy controlling energy is to modify the initial energy (corresponding to regularize the initial condition in terms of dynamic system). This type of methods often focus on modifying our ATNPA’s Augmentation step at Eq. (14).
Stochastic Masking:
Stochastic-mask (random-mask) based models randomly mask or drop edges/nodes of the original graph, corresponding to changes in ATNPA’s Augmentation step at Eq. (14), and then use resulted stochastic graph for propagation. Examples include DropEdge, Drop-connect [43], DropMessage [44].
Combined from the two principles stated from with either boundary condition control and system update control (corresponding to a complete solver for PDE and ODE system), inspired from dynamic system modeling in physics, a branch of method focus on explicit construct continuous ODE or PDE related to graph diffusion and discretize it. As solving ODE or PDE requires both boundary condition (initial condition) and update equation, this type of methods falls into our ATNPA’s Augmentation step at Eq. (14) and Propagation step at Eq. (17).
Physics-inform Process:
@@
Physics-informed methods first model a continuous ODE or PDE related to graph diffusion equation and then discretize the continuous equation with different discretization methods. This leads to nuanced recurrence relation and possibly a combined propagation matrix learnt from both features and topology, which corresponds to ATNPA’s Augmentation step at Eq. (14) and Propagation step at Eq. (17). Examples include GraphCon, GRAND [25], ACMP [46], Neural Sheaf Diffusion [47].
Finally, inspired by recently emerging transformer techniques in sequence generation problems, which can learn long-context sequence relationship, an ensemble of capturing both local relation with shallow layer GNN learning and global relation learning with transformers are studied. It is also worthnoting that transformer structure can be treated as sequence propagating on a dynamic complete graph.
Graph Transformer:
@@
4.2 Residual Connection Methods
@@
Notice that Eq. (20) focus solely on preserve original feature information (skip-connection) without residual connection and no non-linearity function is used for feature update.
@@
where its update equation Eq. (22) combines both residual connection and its skip connection. The residual coefficient remains to be hyperparameter.
@@
where $\alpha+\beta=c_{min}$ and $c_{min}$ is a positive hyperparameter chosen to satisfy the lower bound of initial Dirichlet energy. Its main motivation is to choose an appropriate $c_{min}$ which is an lower bound of initial Dirichlet energy to keep the Dirichlet energy in a controllable range during propagation. Therefore, EGNN is an explicit energy preserving technique compared with implicit energy control by physics-informed methods.
@@
$H^{l}\leftarrow(1-\sigma(\varepsilon_{DE}(F_{\theta}(H^{l-1},A)))H^{l-1}+\sigma(\varepsilon_{DE}(F_{\theta}(H^{l-1},A)))H^{l-1}F_{\theta}(H^{l-1},A)$ (25)
@@
Discussion:
Note that all above methods fit into ATNPA’s unified view by making changes to the Propagation (Eq. (17)) steps. In general, residual based methods have been primarily focused on learning coefficients for each component (i.e., coefficients for $H^{l}$, $H^{l-1}$, or $F_{\theta}(H^{l-1},A)$ etc.). Nevertheless, there is insufficient study and theoretical analysis about the order of each residual components in terms of their position w.r.t activation function $\sigma(\cdot)$. To date, GEN is the only work that empirically validated that order they proposed works better than GCNII.
4.3 Dense Concatenation Methods
@@
It can be observed that the two dense-based methods share similar final aggregation scheme (i.e., final embedding can be considered as a linear combination of layer embeddings). Yet, the embedding learnt in the intermediate layers is different. JKnetcat still includes learnable parameters in the middle and keep non-linearity while DAGNN removes both parts. Without non-linearity and learnable paramters, DAGNN essentially becomes a diffusion kernel based method similar to [57]. We can see that both methods fit into ATNPA’s unified view in Transformation Eq. (15) and Aggregation Eq. (18), which is the key component for dense connection based method.
@@
which is still a variant of ATNPA’s Aggregation (Eq. (15)) with a slight difference in the order of aggregation before convolution instead of after convolution.
Discussion:
@@
4.4 Stochastic-masking Methods
@@
where $\texttt{Bern}(p)$ is a matrix filled with Bernoulli distribution elements and $p$ controls the drop rate. The motivation behind edge dropping is the subspace theorem [19] that indicates less connected graph leading to slow convergence of oversmoothing state. The random-mask modification fits into ATNPA’s Augmentation step Eq. (14).
@@ Discussion:
4.5 Energy Control Methods
Energy-based methods share common motivation of controlling generated embeddings in each layer with constraints on either preserving Dirichlet energy or reducing feature variance. Examples of preserving Dirichlet energy include EGNN, G2-gating, PairNorm, GroupNorm, while NodeNorm reduce feature variance and WeightRep directly reparameterize the learnable parameters to allow independence of input features .
@@ 4.5 Energy Control Methods
WeightRep [45] suggests that learning a proper weight can avoid oversmoothing at any layer and dynamically construct an input-dependent weight and show both empirically and theoretically that such weight can avoid oversmoothing. However, it is observed that performance degradation persists after deep layers, suggesting other factors that might cause the performance degradation.
Discussion:
We note that both normalization and coefficient computation approaches fit into ATNPA’s unified view in Normalization Eq. (16) and Propagation Eq. (17). Direct normalization on features such as mean substraction and variance shifting empirically reduce model capacity and expressive power while coefficients learning seem to be a more promising direction to not only keep Dirichlet energy but also preserve model expressive power.
4.6 Physics-inform Process
Physics-informed methods consider GNN learning as a continuous system and derive solutions by formulating the system’s evolving as a model propagation process. In this context, the time component $t$ in continuous system corresponds to GNN based model’s layer concept. Different discretization methods provide a complex family of methods indicated by a continuous system and most GNN based backbone can be considered as an explicit Euler discretization (only considering the recurrence relation or the Update function in GNN framework) [25]. Examples of continuous systems include GRAND, GraphCon, ACMP, Neural Sheaf Diffusion, and G2-gating (It was first reviewed as GNN backbones, but is also related to the continuous system).
GRAND [25] leverages graph diffusion PDE equations as the continuous system and performs both explicit and implicit Euler discretization. The diffusivity is modeled with an attention structure Eq. (31) related to node features and edges Eq. (32):
@@ 4.6 Physics-inform Process
$U^{\prime}\leftarrow F_{\theta}(A,H,t)-\gamma H-\alpha U$ (33)
$H^{\prime}\leftarrow U$ (34)
where $H^{\prime}$ is the first order derivative with respect to time $t$ (a default setting at physics) and $U^{\prime}$ is equivalent to $H^{\prime\prime}$. After discretization, $t$ is essentially equivalent to layer $l$ in GNN backbones. Following IMEX (implicit-explicit) time discretization [58], GraphCon obtains a new recurrence:
$U^{n}\leftarrow U^{n-1}$ $+\Delta(t)(\sigma(F_{\theta}(A,H^{n-1},t^{n-1}))$
$-\gamma H^{n-1}-\alpha U^{n-1})$ (35)
$H^{n}\leftarrow H^{n-1}$ $+\Delta(t)U^{n}$ (36)
@@
Discussion:
Diffusion systems above share a common point of establishing a connection between the feature changing rate $H^{\prime}$ and the graph gradient $\sum_{j\in Neighbor(i)}|h_{i}-h_{j}|$ which is Dirichlet energy for one node. A discretization of the system then provides a unique complex recurrence relation that preserves established connections.
The niche of diffusion-based methods stem from the design that the system preserves Dirichlet energy (mitigates oversmoothing) through the complex residual recurrence structure, avoiding fixed point convergence at exponential rate and small perturbation deviates the fixed point away in the GraphCon case. This makes diffusion-based method unique, compared with other works that preserve energy through explicit feature value control or coefficient control.